OpenStudy (anonymous):

In a box of candy, there are 7 times as many candies with nuts as candies without nuts. Johnny pulls a candy out of the box at random. What is the probability that Johnny pulled out a candy without nuts?

OpenStudy (kropot72):

Let the number of candies without nuts be n. Then the total number of candies is given by: n + 7n = 8n The probability that Johnny pulled out a candy without nuts is therefore: \[\large P(w/o\ nuts)=\frac{n}{8n}=you\ can\ complete\]

OpenStudy (kropot72):

Hint: Divide the numerator and the denominator by n.

OpenStudy (kropot72):

@Davidsgirl Are you there?

OpenStudy (anonymous):

Im here

OpenStudy (kropot72):

If we divide the numerator by n we get: \[\large \frac{n}{n}=\ ?\]

OpenStudy (anonymous):

okay

OpenStudy (kropot72):

What do you get if you divide n by n? \[\large \frac{n}{n}=can\ you\ answer?\]

OpenStudy (anonymous):

I got it but thank you